OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

slpoe=-1/4 y+3=-1/4(x-2) 4y+12=-x+2 4y+x+10=0

OpenStudy (anonymous):

since the slope of 2 perpendicular lines is slope of L1 x slope of L2 =-1